OpenStudy (anonymous):

y=sqrt of x+5 find intercepts

OpenStudy (asnaseer):

for any curve, the places where it crosses the x- and y-axes are called the intercepts. you also know that y must be zero at the point at which the curve crosses the x-axes. similarly, x must be zero at the point at which the curve crosses the y-axes. therefore, to find the intercepts, you just need to first set y=0 in your equation and solve for x. then set x=0 and solve for y. hope that makes sense?

OpenStudy (anonymous):

i get the process but it the algebra part im not sure about when i set y to zero then do i substract 5 im stuck there

OpenStudy (asnaseer):

so you are given:\[y=\sqrt{x+5}\]is that right? or is it \[y=\sqrt{x}+5\]

OpenStudy (asnaseer):

@SoloN8v88 - which one is the correct equation - the first one I typed or the second one?

OpenStudy (anonymous):

yes the the first one

OpenStudy (asnaseer):

ok, so you are given:\[y=\sqrt{x+5}\]and asked to find the intercepts. lets start with the x-intercept which is where y=0, this gives:\[0=\sqrt{x+5}\]now square both sides to get:\[0^2=x+5\]\[o=x+5\]finally subtract 5 from both sides to get:\[-5=x\]so now you know that the x-intercept is at the point \((-5,0)\) follow the same principal to find the y-intercept.

OpenStudy (anonymous):

oh crap the sq rt of x+is actually being multiplied by x

OpenStudy (anonymous):

x+5

OpenStudy (asnaseer):

you can use the Draw button below - that might make it easier for you to enter your equations

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):